# Explosive synchronization in networks of Type-I neurons with electrical synapses

### v1

@@ Abstract
Explosive synchronization (ES), which was observed in the scale-free network of the Kuramoto model Gómez-Gardeñes et al. (2011), has been studied widely in the oscillator model. However, investigations of ES in neuronal networks, in spite of their importance in neuroscience, are limited and restricted to specific models. In this work, we explore the nature of the transition to synchronization in a class of neurons, namely Type-I neurons. Leveraging the mapping between Type-I neurons and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we investigate whether the conditions known to induce ES in the Kuramoto model also do so in Type-I neurons. The neurons are coupled through electrical synapses and placed on a scale-free and star networks with complete and partial degree-frequency correlation conditions. Our simulations show ES in networks of Quadratic Integrate and Fire (QIF) neurons, the normal form of Type-I neurons, under weak heterogeneity. We further confirm this phenomenon in networks of Morris-Lecar neurons, in the regime of Type-I excitability, under similar conditions to the QIF neurons. Thus, this work establishes a set of universal conditions that allows ES to arise in Type-I neurons.
@@ I Introduction
Despite the significance of ES to better understand brain processes, studies showing ES in neuronal networks are sparse. The neuron models studied exhibit ES under specific network topologies and with additional model-specific constraints. Izhikevich neurons and Chialvo neurons undergo explosive synchronization on a scale-free network without degree frequency correlation. Izhikevich neurons also exhibit such a sharp transition in the ER network Khoshkhou and Montakhab (2018); Roy, Poria, and Hens (2021). Small-world networks of Hodgkin-Huxley and Chialvo neurons also undergo an explosive synchronization Khoshkhou and Montakhab (2020); Boaretto et al. (2019). The FHN model shows ES with nonlocal coupling Verma and Ambika (2022) and on a degree-frequency correlated scale-free network in the presence of noise Chen et al. (2013). The nature of the transition - first order or second order - in neuronal networks can also depend on the type of synapses connecting the neurons Khoshkhou and Montakhab (2018, 2020). The brain consists of a large variety of neurons, each requiring a specific model. The nature of transitions in these neuronal networks is model specific, limiting a broad understanding of synchronization transitions in the Brain network. To overcome this barrier, we explore explosive synchronization in a class of neurons.
@@
For this work we chose Type-I neurons, a class that is characterized by arbitrarily low frequencies near the onset of oscillations. Type-I neurons are predominantly found in the cortex and hippocampus Prescott, De Koninck, and Sejnowski (2008); Laing et al. (2003). Transition from rest to oscillations in these neurons occurs through the Saddle Node on Invariant Circle bifurcation. The Quadratic Integrate-Fire (QIF) model represents the normal form of all Type-I neurons close to the bifurcation point. Taking the QIF model as a prototype of all Type-I neurons, we explore explosive synchronization transition in networks of QIF neurons. To examine the generality of the results obtained, we further analyze synchronization transitions for networks of Morris-Lecar (ML) neurons, known to display Type-I dynamics Tsumoto et al. (2006). This is a biophysical model originally developed to model the dynamics of barnacle muscle fiber Morris and Lecar (1981).
The networks of QIF neurons with weak heterogeneity and weak electrical coupling map to the standard Kuramoto model Clusella, Pietras, and Montbrió (2022). Building on the well documented studies of explosive synchronization in the Kuramoto oscillator network under various conditions, we explore the emergence of explosive synchronization in Type-I neuronal networks. Brain connectivity is known to exhibit a scale-free topologyVan Den Heuvel and Sporns (2013); van den Heuvel and Sporns (2011). Numerous studies with the Kuramoto model on scale free networks have shown explosive Synchronization under different conditions. We leverage the QIF-Kuramoto model mapping to study the synchronization transition of scale-free networks of Type-I neurons with varying degrees of correlation between structure and dynamics.
@@
This paper is organized as follows. In section II we describe the neuron models used in this work with details of the models, commenting on the nature of the bifurcations in these models. In sections III.1 and III.2, results of numerical simulations carried out for networks of QIF neurons and ML neurons are presented, respectively. Section IV gives the concluding remarks of the results discussed in the previous sections.
@@ II Model
In Type-I neurons, the change from excitatory to oscillatory state is achieved by the injected current and occurs through a saddle Node on invariant circle (SNIC) bifurcation. The Quadratic Integrate-and-Fire (QIF) model is a simple yet powerful neuronal model that captures the essential dynamics of Type-I neurons. The following equations describe the QIF model:
$\tau\dot{V_{i}}=V_{i}^{2}+\eta_{i}+\varepsilon I_{i,syn}(t)$ (1a)
$if\ V_{i}>V_{p},\ then\ \ V_{i}\rightarrow V_{r}$ (1b)
where $V_{i}$ is the voltage of $i^{th}$ neuron in the network. The term $\eta_{i}$ represents the external current injected into the neuron, and the term $I_{i,syn}$ represents the net synaptic current flowing into neuron i. When $V_{i}$ reaches a large threshold value, $V_{p}$ (often approximated as $\infty$), it is reset to a large negative value, $V_{r}$ (often taken as -$\infty$), to mimic the action potential firing and reset process of a neuron. Eq. (1a) describes the approach of a QIF neuron to the threshold by the normal form of SNIC bifurcation, while Eq. (1b) represents the reset process. When $\eta_{i}<0$, the neuron is in an excitable state, and if $\eta_{i}>0$, it oscillates, mimicking a train of action potentials (spikes). The frequency of spiking increases gradually with $\eta_{i}>0$ from an arbitrarily low value - a characteristic of Type-I neurons. A well known variable transformed version of the QIF model is the $\Theta$-neuron model" Ermentrout and Kopell (1986). In a recent work Clusella, Pietras, and Montbrió (2022), the QIF model was shown to reduce to the Kuramoto-Sakaguchi model with $V_{p}=-V_{r}=\infty$, under weak synaptic coupling and weak heterogeneity.
In our work, we couple the QIF neurons through electrical synapses, leading to the following form for Eq. (1a).
@@
while the reset condition remains the same. In the above equation, g is the strength of the electrical synapse, $A$ is the adjacency matrix which characterizes the network topology with $A_{ij}=1$ if $i^{th}$ neuron is connected to the $j^{th}$ neuron, and $A_{ij}=0$ if there is no connection between them. There exists no self-connection, i.e., $A_{ii}=0$ for all $i$. We consider two network topologies - scale-free (SF) and star networks. The SF network has a degree distribution given by the power law $p(k)=k^{-\gamma}$ where the degree $k_{i}=\sum_{j}^{N}A_{ij}$ of neuron $i$ is the number of neurons in the network connected to it. In a star network, the single hub neuron is connected to all the other neurons, which are not themselves interconnected. An SF network is characterized by hub-like structures. Hence, star networks act as a precursor to the SF network. We consider networks of QIF neurons with varying degrees of correlation between their degrees and the frequencies. Since for a QIF neuron, $\eta$ is proportional to its frequencyErmentrout and Kopell (1986), degree-frequency correlation is imposed following the rule $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$.
@@
The Morris-Lecar model is a conductance based biophysical model formulated to describe the dynamics of the giant barnacle muscle fibre. The following equations constitute the dynamics of a Morris-Lecar neuron connected to other neurons in the network through electrical synapses.
$C_{M}\frac{dV_{i}}{dt}=$ $-{g_{L}}(V_{i}-V_{l})-g_{Ca}M_{\infty}(V_{i}-V_{Ca})-g_{K}W_{i}(V_{i}-V_{K})$ (5a)
@@
$\frac{dW_{i}}{dt}=$ $\frac{W_{\infty}-W_{i}}{\tau_{W}}$ (5b)
In Eq. (5b), $W_{i}$ represents the recovery variable of the $i^{th}$ neuron. $V_{l}$, $V_{Ca}$, $V_{K}$ represents the equilibrium potentials of the leak, $Ca^{2+}$, and $K^{+}$ currents respectively. $g_{l}$, $g_{Ca}$, and $g_{K}$ are the maximum conductance of the corresponding ionic channels. In our work, we have chosen $V_{l}=-60mV$, $V_{Ca}=120mV$, $V_{K}=-80mV$, $g_{l}=2\ mS/cm^{2}$, $g_{Ca}=2\ mS/cm^{2}$, and $g_{K}=4\ mS/cm^{2}$. The capacitance of the neuron membrane, $C_{M}$ is taken as $20\mu F/Cm^{2}$. $M_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{1})/V_{2}\}}]$ and $W_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{3})/V_{4}}\}]$ are the steady-state activation function of the $Ca^{2+}$ and $K^{+}$ currents respectively and $\tau_{w}(V_{i})=1/[\phi\cosh{\{(V_{i}-V_{3})/2V_{4}\}}]$ is the time constant of the $K^{+}$ activation, with $V_{1}=-1.2mV$, $V_{2}=18.0mV$, $V_{3}=12.0mV$, and $V_{4}=17.4mV$. These parameter values, chosen from Tsumoto et al. (2006), ensure that the neurons are operating in the Type-I regime. External current, $I_{i,ext}$, drives the neurons from an excitable to an oscillatory state through SNIC bifurcation. In an oscillatory state, the free-running frequency of an uncoupled neuron increases with the external current. We consider the same network topologies as those for QIF neurons. Degree-frequency correlation is imposed by fixing $I_{i,ext}=I_{0}+\varepsilon k_{i}$. The value of $I_{0}$ is chosen to keep the ML neuron close to the bifurcation point so that its dynamics can be approximated by the QIF model. The phase of a neuron’s dynamics is expressed by
@@
where $T_{n,i}$ and $T_{n+1,i}$ are the $n^{th}$ and $(n+1)^{th}$ spike times respectively of the $i^{th}$ neuron. These phases are used to calculate the instantaneous order parameter and thereby the global order parameter as described earlier.
@@ III Results
In this section, we discuss the results of numerical simulations carried out for various networks of QIF and ML neurons connected through electrical synapses.
@@ III.1 Networks of QIF neurons
We simulated 1000 QIF neurons arranged on SF networks built via a configuration model. This model gives control over the degree exponent $\gamma$ of the degree distribution $p(k)$. We have considered $2<\gamma<3$ in this work. The dynamics of QIF neurons are governed by Eq. (2) along with the reset rule of Eq. (1b). We fix $V_{p}=-V_{r}=750$ and $\tau=1$ for all neurons. Synaptic strength $g$ is identical between all connected neurons. Simulations are performed into steady states for increasing values of $g$, starting at $g=0$, to a maximum $g$ value. Thereafter, simulations are similarly carried out for decreasing values of $g$. Final steady state voltages for the current $g$ are used as the initial conditions for the next $g$ value. At $g=0$, we start with uniformly distributed voltage values. The order parameter is calculated in the steady state for every value of $g$ in forward and backward runs. We solve Eq. (2) using Euler method with a small step size dtdt ($\approx 10^{-4}$) over the duration $T_{f}=10000$. First $5000$ time units are discarded as transients, and the rest are used for evaluating $R$.
@@
Figure 1: A sharp transition to synchronization accompanied by hysteresis in the order parameter is seen for a scale-free network of 1000 QIF neurons with complete degree-frequency correlation. $\gamma=2.4$ (black line) and $\gamma=2.7$ (red line). For both cases $\bar{\eta}=20$, $\varepsilon=0.01$ with solid curves representing forward simulations and dotted lines representing backward simulations.
@@
A complete degree-frequency correlation is imposed by setting $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ for every neuron. We choose $\bar{\eta}>0$ to place neurons in the oscillatory regime. To maintain weak heterogeneity in $\eta$ values across neurons, we take $\bar{\eta}>>\varepsilon$, with $\bar{\eta}=20$ and $\varepsilon\approx 10^{-2}\in(0.005,0.01)$. With a small $\varepsilon$, the coupling is weak between neurons. Variations in the order parameter while increasing and decreasing values of $g$ are shown in Fig. 1. A sharp transition to synchronization along with hysteresis is clearly seen - indicative of an explosive synchronization. This result is robust to the degree exponent of the scale-free network, as shown in fig (1) for different $\gamma$ values.

### v2

@@ Abstract
Explosive synchronization (ES), which was observed in the scale-free network of the Kuramoto model [J. Gómez-Gardeñes et al.,Phys.Rev.Lett.106,128701(2011)], has been studied widely in the oscillator model. However, investigations of ES in neuronal networks, in spite of their importance in neuroscience, are limited and restricted to specific models. In this work, we explore the nature of the transition to synchronization in a class of neurons, namely Type-I neurons. Leveraging the mapping between networks of weakly heterogeneous Type-I neurons and the Kuramoto model [P.Clusella et al, Chaos 32, 013105 (2022)] under weak coupling, we investigate whether the conditions known to induce ES in the Kuramoto model also do so in Type-I neurons. The neurons are coupled through electrical synapses and placed on a scale-free and star networks with complete and partial degree-frequency correlation conditions. Our simulations show ES in networks of Quadratic Integrate and Fire (QIF) neurons, the normal form of Type-I neurons close to SNIC bifurcation, under weak heterogeneity. We further generalize this phenomenon to networks of Type-I Morris-Lecar neurons, under conditions similar to those of QIF neurons. Thus, this work suggests conditions under which ES can arise in Type-I neurons close to SNIC bifurcation under weak heterogeneity and weak electrical coupling.
@@ I Introduction
Despite the significance of ES to better understand brain processes, studies showing ES in neuronal networks are sparse. The neuron models studied exhibit ES under specific network topologies and with additional model-specific constraints. Izhikevich neurons and Chialvo neurons undergo explosive synchronization on a scale-free network without degree frequency correlation Khoshkhou and Montakhab (2018); Roy, Poria, and Hens (2021). Izhikevich neurons also exhibit such a sharp transition in the ER network. Small-world networks of Hodgkin-Huxley and Chialvo neurons also undergo an explosive synchronization Khoshkhou and Montakhab (2020); Boaretto et al. (2019). The FHN model shows ES with nonlocal coupling Verma and Ambika (2022) and on a degree-frequency correlated scale-free network in the presence of noise Chen et al. (2013). The nature of the transition - first order or second order - in neuronal networks can also depend on the type of synapses connecting the neurons Khoshkhou and Montakhab (2018, 2020). The brain consists of a large variety of neurons, each requiring a specific model. The nature of transitions in these neuronal networks is model specific, limiting a broad understanding of synchronization transitions in the Brain network. To overcome this barrier, we explore explosive synchronization in a class of neurons.
@@
For this work, we chose Type-I neurons, a class that is characterized by arbitrarily low frequencies near the onset of oscillations. Type-I neurons are predominantly found in the cortex and hippocampus Prescott, De Koninck, and Sejnowski (2008); Laing et al. (2003). Transition from rest to oscillations in these neurons occurs through the Saddle Node on Invariant Circle bifurcation. The Quadratic Integrate-Fire (QIF) model represents the normal form of all Type-I neurons close to the bifurcation point. Taking the QIF model as a prototype of all Type-I neurons, we explore explosive synchronization transition in networks of QIF neurons. To examine the generality of the results obtained, we further analyze synchronization transitions for networks of Morris-Lecar (ML) neurons, known to display Type-I dynamics Tsumoto et al. (2006). This is a biophysical model originally developed to model the dynamics of barnacle muscle fiber Morris and Lecar (1981).
Mean firing rate approaches based on dimensionality reduction have been successfully employed to analyze bifurcations and synchronization transitions in heterogeneous networks of QIF neurons Laing (2015); Pietras et al. (2019); Laing (2021); Pietras and Montbrió (2024). While some works Laing (2015); Pietras et al. (2019) report SNIC and subcritical Hopf bifurcations in the mean-field equations, these bifurcations alone do not guarantee explosive synchronization, as bistability—and the associated hysteresis—between synchronous and asynchronous states is also required. The derivation of low-dimensional mean-field equations for QIF populations commonly relies on Lorentzian heterogeneity, which allows for analytical tractability, but also imposes restrictions on the forms of heterogeneity that can be studied. In this work, we instead rely on an alternate analytical formulation of QIF networks. In the formulation, a network of electrically coupled QIF neurons with weak heterogeneity and weak electrical coupling map to the Kuramoto model with zero phase lag Clusella, Pietras, and Montbrió (2022). Building on the well documented studies of explosive synchronization in the Kuramoto oscillator network under various conditions, we explore the emergence of ES in Type-I neuronal networks. We carry out numerical simulations of these neuronal networks with varying degrees of correlation between degree and frequency. In particular, focusing on degree-frequency correlated scale-free and star networks of Kuramoto oscillators, as discussed in Gomez-Gardenes et al. Gómez-Gardeñes et al. (2011), along with partially correlated and uncorrelated scale-free networks studied by Pinto et. al., Pinto and Saa (2015).
@@
This paper is organized as follows. In section II we describe the neuron models used in this work with details of the models, commenting on the nature of the bifurcations in these models. In sections III.1 and III.2, results of numerical simulations carried out for networks of QIF neurons and ML neurons are presented, respectively. Additionally, in section III.1, we numerically validate the mapping between the QIF and the Kuramoto model. Section IV gives the concluding remarks of the results discussed in the previous sections.
@@ II Model
In Type-I neurons, the change from excitatory to oscillatory state is achieved by the injected current and occurs through a saddle Node on invariant circle (SNIC) bifurcation. The Quadratic Integrate-and-Fire (QIF) model is a simple yet powerful neuronal model that captures the essential dynamics of Type-I neurons. The following equations describe the QIF model: {subequations} {align} τ˙V_i = V_i^2 + η_i + εI_i, syn(t)
if V_i > V_p, then V_i →V_r where $V_{i}$ is the voltage of $i^{th}$ neuron in the network. The term $\eta_{i}$ represents the external current injected into the neuron, and the term $I_{i,syn}$ represents the net synaptic current flowing into neuron i. When $V_{i}$ reaches a large threshold value, $V_{p}$ (often approximated as $\infty$), it is reset to a large negative value, $V_{r}$ (often taken as -$\infty$), to mimic the action potential firing and reset process of a neuron. Eq. (II) describes the approach of a QIF neuron to the threshold by the normal form of SNIC bifurcation, while Eq. (II) represents the reset process. When $\eta_{i}<0$, the neuron is in an excitable state, and if $\eta_{i}>0$, it oscillates, mimicking a train of action potentials (spikes). The frequency of spiking increases gradually with $\eta_{i}>0$ from an arbitrarily low value - a characteristic of Type-I neurons. A well known variable transformed version of the QIF model is the $\Theta$-neuron model" Ermentrout and Kopell (1986). This is a popular choice when the system is amenable to Ott-Antonsen ansatz and thereby to dimensionality reduction Chandra et al. (2017); So, Luke, and Barreto (2014). However, when interactions between the neurons are mediated by electrical synapses, synaptic coupling needs to be approximated to avoid singularity Ermentrout (2006) in the $\Theta$-neuron model.
In our work, we couple the QIF neurons through electrical synapses, leading to the following form for Eq. (II).
@@
while the reset condition remains the same. In the above equation, g is the strength of the electrical synapse, $A$ is the adjacency matrix which characterizes the network topology with $A_{ij}=1$ if $i^{th}$ neuron is connected to the $j^{th}$ neuron, and $A_{ij}=0$ if there is no connection between them. There exists no self-connection, i.e., $A_{ii}=0$ for all $i$. We consider two network topologies - scale-free (SF) and star networks. The SF network has a degree distribution given by the power law $p(k)=k^{-\gamma}$ where the degree $k_{i}=\sum_{j}^{N}A_{ij}$ of neuron $i$ is the number of neurons in the network connected to it. In a star network, the single hub neuron is connected to all the other neurons, which are not themselves interconnected. An SF network is characterized by hub-like structures. Hence, star networks act as a precursor to the SF network. We consider networks of QIF neurons with varying degrees of correlation between their degrees and the frequencies.
@@
For a network of electrically connected QIF neurons with $\eta_{i}=\bar{\eta}+\varepsilon\chi_{i}$ taking $\bar{\eta}>0$ and $\chi_{i}>0$, heterogeneity is weak if $\varepsilon\chi_{i}<<\bar{\eta}$. Likewise, $\frac{\varepsilon g}{\bar{\eta}}<<1$ keeps the coupling between neurons weak. In a recent work Clusella, Pietras, and Montbrió (2022) with $V_{p}=-V_{r}=\infty$, and assuming weak heterogeneity and weak coupling, a network of electrically coupled QIF neurons shown to be reduced to the Kuramoto model with zero phase lag Clusella, Pietras, and Montbrió (2022) expressed as
$\dot{\theta_{i}}=\bigg(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon\chi_{i}}{\tau\sqrt{\bar{\eta}}}\bigg)+\frac{\varepsilon g}{\tau}\sum_{j}^{N}\sin(\theta_{j}-\theta_{i})$ (4)
where $\omega_{i}=(\frac{2\sqrt{\bar{\eta}}}{\tau}+\frac{\varepsilon\chi_{i}}{\tau\sqrt{\bar{\eta}}})$ and $\lambda=\varepsilon g/\tau$ are the frequency and the coupling strength respectively of this reduced QIF-Kuramoto model.
The Morris-Lecar model is a conductance based biophysical model formulated to describe the dynamics of the giant barnacle muscle fibre. The following equations constitute the dynamics of a Morris-Lecar neuron connected to other neurons in the network through electrical synapses. {subequations}
$C_{M}\frac{dV_{i}}{dt}=$ $-{g_{L}}(V_{i}-V_{l})-g_{Ca}M_{\infty}(V_{i}-V_{Ca})-g_{K}W_{i}(V_{i}-V_{K})$ (5)
@@
$\frac{dW_{i}}{dt}=$ $\frac{W_{\infty}-W_{i}}{\tau_{W}}$ (6)
In Eq. (6), $W_{i}$ represents the recovery variable of the $i^{th}$ neuron. $V_{l}$, $V_{Ca}$, $V_{K}$ represents the equilibrium potentials of the leak, $Ca^{2+}$, and $K^{+}$ currents respectively. $g_{l}$, $g_{Ca}$, and $g_{K}$ are the maximum conductance of the corresponding ionic channels. In our work, we have chosen $V_{l}=-60mV$, $V_{Ca}=120mV$, $V_{K}=-80mV$, $g_{l}=2\ mS/cm^{2}$, $g_{Ca}=4.0\ mS/cm^{2}$, and $g_{K}=8\ mS/cm^{2}$. The capacitance of the neuron membrane, $C_{M}$ is taken as $20\mu F/Cm^{2}$. $M_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{1})/V_{2}\}}]$ and $W_{\infty}(V_{i})=0.5[1+\tanh{\{(V_{i}-V_{3})/V_{4}}\}]$ are the steady-state activation function of the $Ca^{2+}$ and $K^{+}$ currents respectively and $\tau_{w}(V_{i})=1/[\phi\cosh{\{(V_{i}-V_{3})/2V_{4}\}}]$ is the time constant of the $K^{+}$ activation, with $V_{1}=-1.2mV$, $V_{2}=18.0mV$, $V_{3}=12.0mV$, and $V_{4}=17.4mV$. These parameter values, chosen from Tsumoto et al. (2006), ensure that the neurons are operating in the Type-I regime. External current, $I_{i,ext}$, drives the neurons from an excitable to an oscillatory state through SNIC bifurcation. In an oscillatory state, the free-running frequency of an uncoupled neuron increases with the external current. We consider the same network topologies as those for QIF neurons. Degree-frequency correlation is imposed by fixing $I_{i,ext}=I_{0}+\varepsilon k_{i}$. The value of $I_{0}$ is chosen to keep the ML neuron close to the bifurcation point so that its dynamics can be approximated by the QIF model. Conductance-based models of a neuron, like the Morris-Lecar, do not admit an analytical expression for their phase. Hence, for such a periodically spiking neuron, we compute the phase numerically, which increases linearly between successive spikes as
@@
where $T_{n,i}$ and $T_{n+1,i}$ are the $n^{th}$ and $(n+1)^{th}$ spike times respectively of the $i^{th}$ neuron. The global order parameter computed using these phases ($\phi_{i}$) gives the degree of synchrony in the network.
@@ III Results
In this section, we discuss the results of numerical simulations carried out for electrically coupled networks of QIF and ML neurons with varying magnitudes of correlation between degree and frequency. We begin by assessing the validity of the reduction from the QIF network to the reduced QIF–Kuramoto model.
@@ III.1 Networks of QIF neurons
The dynamics of QIF neurons are governed by Eq. (1) along with the reset rule of Eq. (II). We fix $V_{p}=-V_{r}=750$ and $\tau=1$ for all neurons.
Figure 1: comparison of a degree-frequency correlated star network of 31 QIF neurons and the corresponding reduced QIF-Kuramoto model. The solid lines represent the forward g simulation, while the dotted line represents the backward simulation. For these simulations $\bar{\eta}=20$, $\varepsilon=0.008$.
@@
Synaptic strength $g$ is identical between all connected neurons. Simulations are performed into steady states for increasing values of $g$, starting at $g=0$, to a maximum $g$ value. Thereafter, simulations are similarly carried out for decreasing values of $g$. Final steady state voltages for the current $g$ are used as the initial conditions for the next $g$ value. At $g=0$, we start with uniformly distributed voltage values. The order parameter is calculated in the steady state for every value of $g$ in forward and backward runs. Since for a QIF neuron, $\eta$ is proportional to its frequencyErmentrout and Kopell (1986), degree-frequency correlation is imposed following the rule $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$. We choose $\bar{\eta}>0$ to place neurons in the oscillatory regime. We take $\bar{\eta}=20$ and $\varepsilon\approx 10^{-2}\in(0.005,0.01)$, ensuring weak heterogeneity. In this regime, we numerically assess the validity of mapping the QIF model to the reduced QIF-Kuramoto model. In Clusella, Pietras, and Montbrió (2022), comparisons between the two models were made with zero heterogeneity.
We first carry out simulations of a degree-frequency correlated star network of QIF neurons (Eq. (1)) and the corresponding reduced QIF-Kuramoto network (Eq. (4)). Both the networks are solved using the RK4 note1 method with step size $dt=2.5\times 10^{-4}$ and the synchronization transitions are compared. The phases in Eq. (3) is a good approximation to phase in Eq. (4) with the choice of $V_{p}=-V_{r}=750$ (see Appendix A). Thus, the order parameters calculated from these two are comparable. Simulations are carried out till $T_{f}=50000$, and the order parameter is calculated over the final 30000 time units. In these simulations, when $g$ values are less than a certain value, a small randomness is added Leyva et al. (2025) to the $\eta$ values during the backward run. This is done to slightly perturb the system out of the synchronized branch.
Figure 3: Forward simulations of Star networks of QIF neurons with complete degree-frequency correlation. N is the number of neurons in the network. For these simulations $\bar{\eta}=20$, $\varepsilon=0.008$.
@@
Figure 4: A sharp transition to synchronization accompanied by hysteresis in the order parameter is seen for scale-free networks of 1000 QIF neurons with complete degree-frequency correlation. (a) $\gamma=2.4$ (black line) and $\gamma=2.7$ (red line). (b) average of order parameter for 50 realizations of $\gamma=2.5$ SF network. In all cases $\bar{\eta}=20$, $\varepsilon=0.01$ with solid curves representing forward simulations and dotted lines representing backward simulations.
Fig. 1 shows an abrupt transition in the order parameter accompanied by hysteresis, depicting explosive synchronization in the star network of degree-frequency correlated QIF neurons. The backward transition point of QIF networks shows excellent quantitative agreement with that of the corresponding reduced QIF-Kuramoto model. The forward transition points of the two models lie close to one another but exhibit sensitivity to the $\varepsilon$ values (not included in the figure), which is likely due to finite-size effects.
Further, we study the effect of heterogeneity on the backward transition points of star networks of QIF neurons. The mean-field calculation for a fully degree-frequency correlated star network of Kuramoto oscillators yields the critical coupling strength $\lambda_{c}$ for the backward transition below which a stable synchronized state does not exist Gómez-Gardeñes et al. (2011). The corresponding theoretical $\lambda_{c}$ for the star network of QIF neurons is calculated in Appendix B. This gives
For $\varepsilon=0.0095$ and $\varepsilon=0.0065$, theoretically predicted $\lambda_{c}$ values are $\approx 0.0019$ and $\approx 0.0013$ respectively. In Fig. 2, it is seen that the backward transitions occur at $\lambda$’s close to these values predicted by the above equation. The above results validated the mapping between the Kuramoto model and the weakly heterogeneous QIF model with weak electrical coupling. We also performed simulations of QIF star networks of different sizes. Fig. 3 shows forward simulation results. There is a sharp transition to synchronization across different sizes of star networks. Each of these is accompanied by hysteresis (not shown in the figure).
@@
Further, we examine other QIF network conditions that are known to be sufficient to induce ES in the Kuramoto model. In particular, scale-free networks of Kuramoto oscillators with varying degrees of correlation between degree and frequency.
Figure 6: As the heterogeneity ($\varepsilon$) is increased, hysteresis width decreases for the SF networks of 1000 degree-frequency correlated QIF neurons and the transitions become less sharp. In the figure, the solid lines represent the forward simulation, and the dotted lines represent the backward simulation. All the other parameters are considered the same as those taken in Fig. 4a, $\gamma=2.5$ case.
We first simulate 1000 fully degree-frequency correlated QIF neurons arranged on SF networks. The SF networks are built via a configuration model. This model gives control over the degree exponent $\gamma$ of the degree distribution $p(k)$. We have considered $2<\gamma<3$ in this work. For the degree-frequency correlated scale-free networks in our simulations, we take $\bar{\eta}=20$ and $\varepsilon=0.01$. With these parameters, $99\%$ of the neurons have $20.0\leq\eta_{i}\leq 20.5$, maintaining the weak heterogeneity. We solve Eq. (1) using Euler method with a small step size dtdt ($\approx 10^{-4}$) over the duration $T_{f}=10000$. The first $5000$ time units are discarded as transients, and the rest are used for evaluating $R$. Simulations are performed by varying g following the same method as described for star networks. Variations in the order parameter while increasing and decreasing values of $g$ are shown in Fig. 4. A sharp transition to synchronization, along with hysteresis, is clearly seen - exhibiting an explosive synchronization. This result is robust to the degree exponent of the scale-free network, as shown in Fig. 4 for different $\gamma$ values. We carried out simulations over 50 realizations of the SF network with the same $\gamma$. Hysteresis and sharp transitions persist in all of them. Forward transition point is found to be at $g_{fc}$ $0.766\pm 0.092$ and the backward transition point at $g_{bc}=0.453\pm 0.027$. Fig. 4b shows the averaged values of $R$ over these 50 realizations. The Kuramoto model with zero phase lag is known to undergo explosive synchronization even with partial degree-frequency correlation Pinto and Saa (2015). Exploiting the mapping between the QIF neuronal network and the Kuramoto model Clusella, Pietras, and Montbrió (2022), we also investigated synchronization transition in partially degree-frequency correlated QIF networks. To achieve partial degree-frequency correlation in the QIF network, a fraction of the neuron population is assigned frequencies following the $\eta_{i}=\bar{\eta}+\varepsilon k_{i}$ rule. The rest are assigned $\chi_{i}$ randomly from a Lorentzian distribution, making the frequencies random.
[…]
Figure 8: Plots of 1000 ML neurons connected through an SF network. In all cases, the same SF network with $\gamma=2.4$ is used. (a) Complete degree-frequency correlation. (b) Only the top $10\%$ of neurons with the highest degree are assigned frequencies correlated with their degree, while the rest are assigned random frequencies. (c) All neurons are assigned random frequencies. In both (b) and (c), random frequencies are drawn from a Lorentzian distribution $p(k)=a/\pi((k-k_{0})^{2}-a^{2})$ with $k_{0}=0$ and $a=1$. For all cases $I_{0}=43.0\mu A/cm^{2}$ and $\varepsilon=0.01$.